OpenStudy (anonymous):

sint=-3/7, cost=-2 squareroot 10. find tan t

OpenStudy (anonymous):

well, tan(t) = sin(t)/cos(t) so all you have to do is plug in :)

OpenStudy (anonymous):

sint=-3/7, cost=\[2\sqrt{10}\div7 \]. find tan t

OpenStudy (anonymous):

is it 3squarroot10 /20

OpenStudy (anonymous):

sint=-3/7 and cost=2sqrt(10)/7 tant=sint/cost=[-3/7]/[2sqrt(10)/7]=(-3)(7)/[(7)2sqrt(10)]=-3/[2sqrt(10)]